How do you find degree y of a right triangle whose hypotenuse is n (square root 5), adjacent is 2n and opposite side is n.
Answer by Theo(13342)

your angle is 26.56505118 degrees.

you get sin(y) = n / n*sqrt(5) which is equal to 1/sqrt(5)

you get cos(y) = 2n / n*sqrt(5) which is equal to 2/sqrt(5)

you get tan(y) = n / 2n = 1/2

use your calculator to get arcsin and arccos and arctan and they all point to 26.56505118 degrees.